# Ballistic electron transport described by a generalized Schrödinger equation

### v1

@@ Abstract
We propose a Schrödinger equation of arbitrary order for modeling charge transport in semiconductors operating in the ballistic regime. This formulation incorporates non-parabolic effects through the Kane dispersion relation, thereby extending beyond the conventional effective mass approximation. Building upon the framework introduced in G. E. Aliffi, G. Nastasi, V. Romano, ZAMP 76, 155 (2025), we derive a hierarchy of models, each governed by a Schrödinger equation of increasing order. As in the standard second-order case, the problem is formulated on a finite spatial domain with suitable transparent boundary conditions. These conditions are designed to simulate charge transport in a quantum coupler where an active region – representing the electron device – is connected to leads acting as reservoirs. We investigate several analytical properties of the proposed models and derive a generalized expression for the current, valid for any order. This formula includes additional terms that account for interference effects arising from the richer wave structure inherent in higher-order Schrödinger equations, which are absent in the effective mass approximation. Numerical simulations of a resonant tunneling diode (RTD) illustrate the key features of the solutions and highlight the impact of the generalized formulation on device behavior.
@@ 1 Introduction
In [23], a correction of order $\hbar^{4}$ was introduced into the dispersion relation, leading to a fourth-order Schrödinger equation (SE) whose analytical properties were thoroughly investigated. Simulations involving single and double potential barriers with an applied bias revealed interference effects absent in the second-order SE. In this work, we generalize that approach to arbitrary order in the expansion of the Kane dispersion relation. We derive a continuity equation for the single-electron probability density and a general expression for the probability current valid at any order. Transparent boundary conditions are also formulated for higher-order models. We investigate several properties of the proposed framework and present numerical examples to illustrate its behavior. The structure of the paper is as follows. In Section 2, we derive the generalized Schrödinger equation and discuss the functional setting. Section 3 is devoted to the derivation of a general expression for the probability current. In Section 4, we formulate transparent boundary conditions for arbitrary order, enabling the reduction of the problem to a compact domain. Section 6 compares the probability current for the second- and fourth-order SE. Finally, Section 7 presents numerical simulations of an RTD to highlight key features of the proposed models.
@@ 2 Dispersion relation and generalized Schrödinger equation
where $\epsilon(k)$ denotes the electron energy, $k$ is the magnitude of the electron wave vector, and $m^{*}$ is the effective electron mass. For example, in GaAs, $m^{*}=0.067m_{e}$, with $m_{e}$ being the free electron mass. The positive parameter $\alpha$ is referred to as the non-parabolicity factor. In the limit $\alpha\to 0^{+}$, the standard parabolic band approximation is recovered:
@@
where
@@
we obtain the operator form of the energy:
@@
Let $\hat{\Psi}(k,t)$ denote the wave function in wave-vector space. It is related to $\Psi(x,t)$ via the unitary Fourier transform
@@
Equation (2.10) allows for the inclusion of the full dispersion relation, but the potential term becomes integral in nature. However, for the simulation of electron devices such as RTDs, following [8], we aim to solve the Schrödinger equation in a bounded domain representing the active region of the device, with appropriate boundary conditions. Formulating such boundary conditions in wave-vector space is not straightforward; therefore, we focus exclusively on the position-space formulation given by Equation (2.7).
@@ 3 Generalization of the Current
Altogether, we have established the following result:
@@ 4 Transparent boundary conditions for the generalized stationary Schrödinger equation
$H=\sum_{j=1}^{+\infty}c_{j}\alpha^{j-1}\left(\left(-\frac{\hbar^{2}}{2m^{*}}\right)^{j}\Delta^{j}\right)-qV(x)=0.$ (4.1)
@@ 4.1 Case $k_{1}>0$
If we have the coincidence of some wave-vectors, the analysis is much more involved and the situation must be handled on a case-by-case basis.
Now we pass to evaluate the wave-vectors $k_{i}$. In the region $x<0$, after the substitution the ansatz in the Schrödinger equation and exploiting the independence of the functions $e^{-ik_{j}x},\ j=1,\ldots,s$ we get for the wave-vectors ${k}_{j}$ the relation
$\sum_{p=1}^{s}c_{p}\alpha^{p-1}\left(-\frac{\hbar^{2}}{2m^{*}}\right)^{p}(-ik_{j})^{2p}-(qV(0)+E)=0,\ j=1,\ldots,s$ (4.10)
from which it follows that
$E=\sum_{p=1}^{s}c_{p}\alpha^{p-1}\left(-\frac{\hbar^{2}}{2m^{*}}\right)^{p}(ik_{1})^{2p}-qV(0).$ (4.11)
On the other hand, considering the region $x>L$, we get for the wave-vectors $\tilde{k}_{j}$ the relation
@@ 4.2 Case $k_{1}<0$
Electron waves are injected at the boundary point $x=L$, where they may either be reflected back into Region III or transmitted into Region I. The wave function is modeled by the following ansatz
@@
For $x>L$, the wave vectors $k_{j}$ satisfy
@@
If we have the coincidence of some wave-vectors, as for the case $k_{1}>0$ the analysis is much more involved and the situation must be handled on a case-by-case basis.
@@
Only wave vectors with positive real parts are included in the ansatz. If any wave vector is complex, to ensure bounded solutions, we take only the wave vectors having negaitive coefficient of the imaginary part and evanescent modes arise.
@@ 5 Well posedness of the higher order SE with transparent boundary conditions
We establish a well-posedness result for the generalized Schrödinger model of arbitrary even order, thereby extending the result previously proven for the fourth-order case. For clarity, we present the case where $k_{1}>0$; the case $k_{1}<0$ can be treated analogously.
@@
Under the hypothesis on the potential, the coefficients of the linear equations (5.1) are in $L^{1}_{loc}(0,L)$ ensuring the applicability of the existence and uniqueness theorem for the associated Cauchy problem. Let $(\varphi_{0},\varphi_{1},\varphi_{2},\ldots,\varphi_{2s-1})$ be a fundamental set of solutions to 5.1, each satisfying the initial conditions:
@@
The general integral of (5.1) can be expressed as $\Psi(x)=\sum_{p=0}^{2s-1}c_{p}\Psi_{p}(x)$ with $c_{p}\in\mathbb{C}$.
@@ 6 Comparison of the current between the second and the fourth order SE
In Fig.1 we have plotted the ensemble incident, reflected and transmitted probability current, obtained numerically, versus the wave-vector of the incident wave.
Figure 1: Numerical evaluation of the ensemble incident, reflected and transmitted probability current density versus the wave-vector of the incident wave.
@@
Evanescent modes are absent when all wave vectors $k_{i}\in\mathbb{R}$ for any $i=1,2,3,4$. These wave vectors satisfy the quartic equation
$ax^{4}+\frac{\hbar^{2}}{2m^{*}}x^{2}-(qV(0)+E)=(x^{2}-k_{1}^{2})(ax^{2}+\frac{\hbar^{2}}{2m^{*}}+ak_{1}^{2})=0$ (6.11)
@@
$E=\frac{\hbar^{2}}{2m^{*}}k_{1}^{2}+ak_{1}^{4}-qV(0)$ (6.12)
and $a=-\frac{\hbar^{4}}{4(m^{*})^{2}}\alpha<0.$
Equation 6.11 admits only real solutions if and only if $\frac{\hbar^{2}}{2m^{*}}+ak_{1}^{2}>0$, that is, if $\sqrt{-\frac{\hbar^{2}}{2m^{*}a}}<k_{1}<\sqrt{-\frac{\hbar^{2}}{2m^{*}a}}$.
@@ 7 Simulation of a resonant tunneling diode
To evaluate the key quantities relevant for RTD design—namely, the average electron density and current—we consider an incident plane wave with wave vector $k$. The associated density operator in coordinate representation has kernel $\rho_{k}(x,y)=\Psi_{k}(x)\overline{\Psi_{k}(y)}$, where $\Psi_{k}(x)$ is the solution of the corresponding Schrödinger equation. The kernel of the ensemble density matrix operator is given by
@@
$f_{FD}(k)=\dfrac{1}{1+\exp\left(\frac{\epsilon(k)-E_{F}}{k_{B}T}\right)},$
with $E_{F}$ being the Fermi energy and $T$ the absolute temperature which we set 300 $K$ (room temperature). The electron density $n(x)$ is then obtained as
@@
where tr denotes the trace.
Figure 2: RTD type potential in the case $_{0}=0 V, V$$_{L}=0.1 V$, $V_{b}=-0.3$ V, $L=135$ nm.
The total electric current within the device is given by
${\cal J}=-q\frac{g_{s}g_{v}}{(2\pi)^{3}}\int_{\mathbb{R}^{3}}f_{FD}(k)J_{k_{x}}d^{3}k$
By adopting cylindrical coordinates: $k_{x}=k_{x}$, $k_{y}=\sigma\cos\theta$, $k_{z}=\sigma\sin\theta$, with $\sigma\in[0,+\infty[$ and $\theta\in[0,2\pi[$, the expression becomes
${\cal J}=-q\frac{g_{s}g_{v}}{(2\pi)^{2}}\int_{-\infty}^{+\infty}dk_{x}\int_{0}^{+\infty}\frac{J_{k_{x}}\sigma}{1+exp\left(\dfrac{\epsilon(\sigma,k_{x})-E_{F}}{k_{B}T}\right)}d\sigma$ (7.3)
@@
$g(\sigma,k_{x})=\sigma\left[1+\exp\left(\dfrac{\epsilon(\sigma,k_{x})-E_{F}}{k_{B}T}\right)\right]^{-1}$
which numerically exhibits compact support.
@@
Using Gaussian quadrature for numerical integration and solving the Schrödinger equation at each quadrature node with the scheme proposed in [23], we obtain the results shown in Figures 4 and 5. The current is conserved with high numerical accuracy (fluctuations are below $10^{-8}$), confirming the robustness of the computational method. While SE2 and SE4 yield qualitatively similar results, quantitative differences are evident: SE2 produces a higher peak in the resonant region for the electron density, whereas SE4 reveals interference effects. Regarding the average electron current, SE4 yields a value that is approximately 38% of that predicted by SE2.
Figure 4: Electron density obtained with the fourth-order SE (left) and the second order SE (right) in the case $_{0}=0 V, V$$_{L}=0.1 V$, $V_{b}=-0.3$ V, $L=135$ nm by using the Kane dispersion relation.
Figure 5: Electron current density obtained with the fourth-order SE (left) and the second order SE (right) in the case $_{0}=0 V, V$$_{L}=0.1 V$, $V_{b}=-0.3$ V, $L=135$ nm by using the Kane dispersion relation.
For completeness, we also compare results obtained using a parabolic band approximation in the Fermi-Dirac statistics. The corresponding solutions are shown in Figures 6 and 7. The SE4 density exhibits more pronounced oscillations due to interference effects compared to SE2, and its magnitude is higher than in the Kane model. As in the semiclassical case, the parabolic band approximation tends to overestimate the electron current density. SE2 again predicts a higher current than SE4, reinforcing the importance of employing accurate dispersion relations.
@@
Figure 7: Electron current density in the case $_{0}=0 V, V$$_{L}=0.1 V$, $V_{b}=-0.3$ V, $L=135$ nm in the parabolic band approximation. SE4 left, SE2 right.
Conclusion and acknowledgements

### v2

@@ Abstract
We propose a Schrödinger equation of arbitrary order for modeling charge transport in semiconductors operating in the ballistic regime. This formulation incorporates non-parabolic effects through the Kane dispersion relation, thereby extending beyond the conventional effective mass approximation. Building upon the framework introduced in G.E. Aliffi, G. Nastasi, V. Romano, ZAMP 76, 155 (2025), we derive a hierarchy of models, each governed by a Schrödinger equation of increasing order. As in the standard second-order formulation, the problem is formulated on a bounded spatial domain with suitable transparent boundary conditions. These conditions are designed to simulate charge transport in a quantum coupler where an active region – representing the electron device – is connected to leads acting as reservoirs. We investigate several analytical properties of the proposed models and derive a generalized expression for the current, valid for any order. This formula includes additional terms that account for interference effects arising from the richer wave structure inherent in higher-order Schrödinger equations, which are not captured by the effective mass approximation. Numerical simulations of a resonant tunneling diode (RTD) illustrate the key features of the solutions and highlight the impact of the generalized formulation on device behaviour.
@@
Author Contribution declaration
Giovanni Nastasi: Formal analysis, Investigation, Methodology, Supervision, Validation, Visualization, Writing – original draft, Writing .
Giulia Aliffi: Formal analysis, Investigation, Methodology, Validation, Visualization, Writing – original draft, Writing .
Vittorio Romano: Conceptualization, Formal analysis, Funding acquisition, Investigation, Methodology, Supervision, Validation, Writing – original draft, Writing .
@@ 1 Introduction
In [23], a correction of order $\hbar^{4}$ was introduced into the dispersion relation, leading to a fourth-order Schrödinger equation (SE) whose analytical properties were thoroughly investigated. Simulations involving single and double potential barriers with an applied bias revealed interference effects absent in the second-order SE. In this work, we generalize that approach to arbitrary order in the expansion of the Kane dispersion relation. We derive a continuity equation for the single-electron probability density and a general expression for the probability current valid at any order. Transparent boundary conditions are also formulated for higher-order models. We investigate several properties of the proposed framework and present numerical examples to illustrate its behavior. The structure of the paper is as follows. In Section 2, we derive the generalized Schrödinger equation and discuss the functional setting. Section 3 is devoted to the derivation of a general expression for the probability current. In Section 4, we formulate transparent boundary conditions for arbitrary order, enabling the reduction of the problem to a compact domain. Section 5 is dedicated to the well-posedness of the resulting boundary value problem. Section 6 compares the probability current for the second- and fourth-order SE. Finally, Section 7 presents numerical simulations of an RTD to highlight key features of the proposed models.
@@ 2 Dispersion relation and generalized Schrödinger equation
where $\epsilon(k)$ denotes the electron energy measured from the bottom of the conduction band $E_{C}$, $k$ is the magnitude of the electron wave vector, and $m^{*}$ is the effective electron mass. For example, in GaAs, $m^{*}=0.067m_{e}$, with $m_{e}$ being the free electron mass. The positive parameter $\alpha$ is referred to as the non-parabolicity factor. In the limit $\alpha\to 0^{+}$, the standard parabolic band approximation is recovered:
@@
where $\binom{\beta}{j}$ denotes the generalized binomial coefficient, i.e.
@@
we obtain the operatorial form of the energy:
@@
Let $\hat{\Psi}(k,t)$ denote the wave function in wave vector space. It is related to $\Psi(x,t)$ via the unitary Fourier transform
@@
Equation (2.10) allows for the inclusion of the full dispersion relation, but the potential term becomes integral in nature. However, for the simulation of electron devices such as RTDs, following [8], we aim to solve the Schrödinger equation in a bounded domain representing the active region of the device, with appropriate boundary conditions. Formulating such boundary conditions in wave vector space is not straightforward; therefore, we focus exclusively on the position-space formulation given by Equation (2.7).
@@ 3 Generalization of the Current
Altogether, we have established the following result.
@@ 4 Transparent boundary conditions for the generalized stationary Schrödinger equation
$H=\sum_{j=1}^{+\infty}c_{j}\alpha^{j-1}\left(\left(-\frac{\hbar^{2}}{2m^{*}}\right)^{j}\Delta^{j}\right)-qV(x).$ (4.1)
@@ 4.1 Case $k_{1}>0$
If we have the coincidence of some wave vectors, the analysis is much more involved and the situation must be handled on a case-by-case basis.
Now we pass to evaluate the wave vectors $k_{i}$. In the region $x<0$, after the substitution the ansatz in the Schrödinger equation and exploiting the independence of the functions $e^{-ik_{j}x},\ j=1,\ldots,s$ we get for the wave vectors ${k}_{j}$ the relation
$\sum_{p=1}^{s}c_{p}\alpha^{p-1}\left(-\frac{\hbar^{2}}{2m^{*}}\right)^{p}(-ik_{j})^{2p}-(qV(0)+E)=0,\ j=1,\ldots,s.$ (4.10)
By using the fact that $k_{1}$ is known, from the previous relation for $j=1$ we evaluate the energy $E$
$E=\sum_{p=1}^{s}c_{p}\alpha^{p-1}\left(-\frac{\hbar^{2}}{2m^{*}}\right)^{p}(ik_{1})^{2p}-qV(0)$ (4.11)
while the remaining relations allow us to get the wave vectors $k_{j}$, $j=2,\ldots,s$.
On the other hand, considering the region $x>L$, we get for the wave vectors $\tilde{k}_{j}$ the relation
@@ 4.2 Case $k_{1}<0$
Electron waves are injected at the boundary point $x=L$, where they may either be reflected back into Region III or transmitted into Region I. The wave function is modelled by the following ansatz
@@
while the remaining relations gives the values of the remaining $k_{j}$’s, $j=2,\ldots,s$. For $x>L$, the wave vectors $k_{j}$ satisfy
@@
If we have the coincidence of some wave vectors, as for the case $k_{1}>0$ the analysis is much more involved and the situation must be handled on a case-by-case basis.
@@
Only wave vectors with positive real parts are included in the ansatz. If any wave vector is complex, to ensure bounded solutions, we take only the wave vectors having negative coefficient of the imaginary part and evanescent modes arise.
@@ 5 Well posedness of the higher order SE with transparent boundary conditions
We establish a well-posedness result for the generalized Schrödinger model of arbitrary even order, thereby extending the result previously proven for the fourth-order case. For the sake of clarity, we present the case where $k_{1}>0$; the case $k_{1}<0$ can be treated analogously.
@@
Under the hypothesis on the potential, the coefficients of the linear equations (5.1) are in $L^{1}_{loc}(0,L)$ ensuring the applicability of the existence and uniqueness theorem for the associated Cauchy problem. Let $(\varphi_{0},\varphi_{1},\varphi_{2},\ldots,\varphi_{2s-1})$ be a fundamental set of solutions to 5.1, satisfying the initial conditions:
@@
The general integral of (5.1) can be expressed as $\Psi(x)=\sum_{p=0}^{2s-1}c_{p}\varphi_{p}(x)$ with $c_{p}\in\mathbb{C}$.
@@ 6 Comparison of the current between the second and the fourth order SE
In Fig.1 we have plotted the ensemble incident, reflected and transmitted probability current, obtained numerically, versus the wave vector of the incident wave.
Figure 1: Numerical evaluation of the ensemble incident, reflected and transmitted probability current density versus the wave vector of the incident wave.
@@
Evanescent modes are absent when all wave vectors $k_{i}\in\mathbb{R}$ for any $i=1,2,3,4$. The wave vectors $k_{1}$ and $k_{2}$ satisfy the quartic equation
$ax^{4}+\frac{\hbar^{2}}{2m^{*}}x^{2}-(qV(x_{b})+E)=(x^{2}-k_{1}^{2})(ax^{2}+\frac{\hbar^{2}}{2m^{*}}+ak_{1}^{2})=0$ (6.11)
@@
$E=\frac{\hbar^{2}}{2m^{*}}k_{1}^{2}+ak_{1}^{4}-qV(x_{b}),$ (6.12)
$a=-\frac{\hbar^{4}}{4(m^{*})^{2}}\alpha<0$ and $x_{b}=0$ if the electron enters from the left contact, $x_{b}=L$ if the electron enters from the right contact. Equation 6.11 admits only real solutions if and only if $\frac{\hbar^{2}}{2m^{*}}+ak_{1}^{2}>0$, that is, if $\sqrt{\dfrac{2m^{*}}{\alpha\hbar^{2}}}<k_{1}<\sqrt{\dfrac{2m^{*}}{\alpha\hbar^{2}}}$.
The other wave vectors are given by
$k_{3,4}=\sqrt{-\frac{\hbar^{2}}{4m^{*}a}\pm\frac{1}{2a}\sqrt{\frac{\hbar^{4}}{4(m^{*})^{2}}+4a(qV(x_{c})+E)}}$
where in the complex case the root must be intended with positive real and imaginary parts. Here $x_{c}=0$ if the electron enters from the right contact, $x_{b}=L$ if the electron enters from the left contact. Therefore a priori it is not possible to exclude that $k_{3}$ or $k_{4}$ lead to evanescent waves.
@@ 7 Simulation of a resonant tunneling diode
Figure 2: RTD type potential energy in the case $_{0}=0 V, V$$_{L}=0.1 V$, $V_{b}=-0.3$ V, $L=135$ nm.
Figure 3: Plot of the full Kane dispersion relation $\epsilon_{K}$ (continuous line), fourth order approximation $\epsilon_{4}$ (dashed line) and second order approximation $\epsilon_{2}$ (dotted line) along with the Fermi-Dirac distribution evaluated with the previous expressions of the energy band, denoted as $f_{FD_{K}}$, $f_{FD_{4}}$, $f_{FD_{2}}$, respectively. When the quartic approximation is adopted spurious tails are present close to $\pm\sqrt{\frac{2m^{*}}{\alpha\hbar^{2}}}$ but they are well outside from the region of physical interest.
To evaluate the key quantities relevant for RTD design - namely, the average electron density and current - we consider an incident plane wave with wave vector $k$. The associated density operator in coordinate representation has kernel $\rho_{k}(x,y)=\Psi_{k}(x)\overline{\Psi_{k}(y)}$, where $\Psi_{k}(x)$ is the solution of the corresponding Schrödinger equation. The kernel of the ensemble density matrix operator is given by
@@
$f_{FD}(k)=\dfrac{1}{1+\exp\left(\frac{E(k)-E_{F}}{k_{B}T}\right)},$
with $E(k)$ being the electron energy, $E_{F}$ the Fermi energy and $T$ the absolute temperature which we set 300 $K$ (room temperature). We observe that
$E(k)-E_{F}=E(k)-E_{C}-(E_{F}-E_{C})=\epsilon(k)-\mu$
[…]
In principle the Kane approximation is considered valid for any wave vector in $\mathbb{R}^{3}$. In Fig. 3 we have plotted $\epsilon(k)$ along with $f_{FD}(k)$ versus the modulus $k$ of the wave vector comparing the full Kane expression with the quartic and quadratic (parabolic) approximation. One observes that from a numerical point of view $f_{FD}(k)$ exhibits a compact support. Within such a set the full Kane dispersion relation is accurately approximated by the quartic expansion; indeed, also the quadratic approximation is acceptable, although slightly less precise. Fig. 3 further shows that the quartic approximation is no longer monotone for $k>\sqrt{\dfrac{m^{*}}{\alpha\hbar^{2}}}$. The loss of monotonicity is not problematic since the interval $\left[-\sqrt{\dfrac{m^{*}}{\alpha\hbar^{2}}},\sqrt{\dfrac{m^{*}}{\alpha\hbar^{2}}}\right]$ fully contains the support of $f_{FD}(k)$. Spurious tails of the distribution function with the quartic approximation appear close to $\pm\sqrt{\dfrac{2m^{*}}{\alpha\hbar^{2}}}$ but these lie outside the domain of numerical integration. It is worth noting that, despite the similarity of the resulting statistics, the parabolic and quartic dispersion relation leads to different Schrödinger equations. So one expects different results for the average quantities of interest.
The electron density $n(x)$ is obtained as
@@
where tr denotes the trace. The total electric current density is given by
${\cal J}=-q\frac{g_{s}g_{v}}{(2\pi)^{3}}\int_{\mathbb{R}^{3}}f_{FD}(k)J_{k_{x}}d^{3}k.$
By adopting cylindrical coordinates: $k_{x}=k_{x}$, $k_{y}=\sigma\cos\theta$, $k_{z}=\sigma\sin\theta$, with $\sigma\in[0,+\infty[$ and $\theta\in[0,2\pi[$, the latter becomes
${\cal J}=-q\frac{g_{s}g_{v}}{(2\pi)^{2}}\int_{-\infty}^{+\infty}dk_{x}\int_{0}^{+\infty}\frac{J_{k_{x}}\sigma}{1+\exp\left(\dfrac{\epsilon(\sigma,k_{x})-\mu}{k_{B}T}\right)}d\sigma,$ (7.3)
@@
$g(\sigma,k_{x})=\sigma\left[1+\exp\left(\dfrac{\epsilon(\sigma,k_{x})-\mu}{k_{B}T}\right)\right]^{-1}.$
which numerically exhibits a compact support consistently with the consideration outlined above. Therefore we restrict the numerical domain to $\overline{k}\leq k_{x}\leq\overline{k}$ and $0\leq\sigma\leq\overline{\sigma}$, with $\overline{k}=$1.5 nm-1 and $\overline{\sigma}=$ 1 nm-1. Within this range the full Kane’s dispersion relation and the quartic approximation are in good agreement as shown by Figure 5 which in logarithmic scale highlights in detail the differences between $f_{FD}(k)$ with the full Kane dispersion relation and with the quartic and quadratic approximation.
@@
Figure 5: Logarithmic plot of $f_{FD}(k)$ and the energy band with the full Kane dispersion relation and with the quartic and quadratic approximation. The notation is as in Fig. 3
@@
Using Gaussian quadrature for numerical integration and solving the Schrödinger equation at each quadrature node with the scheme proposed in [23], we obtain the results shown in Figures 6 and 7. Regarding the density while SE2 and SE4 yield qualitatively similar results, quantitative differences are evident: SE2 produces a higher peak in the resonant region for the electron density, whereas SE4 reveals interference effects. The density obtained with SE4 has only a slight difference between the case of the full Kane dispersion and the case of the quartic approximation.
[…]
Figure 6: Electron density obtained with the fourth-order SE (continuous line, left) , and the second order SE (right) in the case $_{0}=0 V, V$$_{L}=0.1 V$, $V_{b}=-0.3$ V, $L=135$ nm by using the Kane dispersion relation. For comparison also the electron density with the quartic approximation in the statistics (dashed line, left) is reported
Figure 7: Electron current density obtained with the fourth-order SE by using the Kane dispersion relation and the second order SE in the case $_{0}=0 V, V$$_{L}=0.1 V$, $V_{b}=-0.3$ V, $L=135$ nm by using the Kane dispersion relation. For comparison also the electron current density with the quartic approximation in the statistics is reported.
For completeness, we also compare results obtained using a parabolic band approximation in the Fermi-Dirac statistics. The corresponding solutions are shown in Figures 8 and 9. The SE4 density exhibits more pronounced oscillations due to interference effects compared to SE2, and its magnitude is higher than in the Kane model. As in the semiclassical case, the parabolic band approximation tends to overestimate the electron current density in absolute value. SE2 again predicts a higher current than SE4, reinforcing the importance of employing accurate dispersion relations.
@@
Figure 9: Electron current density in the case $_{0}=0 V, V$$_{L}=0.1 V$, $V_{b}=-0.3$ V, $L=135$ nm in the parabolic band approximation. SE4: continuous line. SE2: dashed line.
Conclusion